# MotionStreamer: Streaming Motion Generation via Diffusion-based Autoregressive Model in Causal Latent Space

### v2

@@ A Implementation Details
For the Causal TAE, both the encoder and decoder are based on the 1D causal ResNet blocks [20]. The temporal downsampling rate $l$ is set to 4 and all motion sequences are cropped to $N=64$ frames during training. We train the first 1900K iterations with a learning rate of 5e-5 and the remaining 100K iterations with a learning rate of 2.5e-6. We use the AdamW optimizer [32] with $[\beta_{1},\beta_{2}]=[0.9,0.99]$ and a batch size of 128. We provide an ablation study on the hyperparameter $\lambda$ of root loss $L_{root}$ in Tab. 5. The latent dimension $d_{c}$ and hidden size are set to 16 and 1024, respectively. The latent dimension significantly impacts the compression rate, while the hidden size affects the model’s capacity. Both factors influence reconstruction and subsequent generation quality, requiring a careful trade-off between compression efficiency and generative performance. Ablation studies on the latent dimension and hidden size are provided in Tab. 6. To further improve the quality of the reconstructed motion, we add a linear layer after the embedded Gaussian distribution parameters as a latent adapter to get a lower-dimensional and more compact latent space for subsequent sampling, as proposed in [11].
For the Transformer inside the AR model, we use the architecture akin to LLaMA [55] with 12 layers, 12 attention heads and 768 hidden dimension. The ablation for different scales of the Transformer is provided in Tab. 7. Block size is set to 78 and we choose RoPE [52] as the positional encoding. For the diffusion head after Transformer, we use MLPs with 1768 hidden dimension and 9 layers. The output vectors of the Transformer serve as the condition of denoising via AdaLN [38]. We adopt a cosine noise schedule with 50 steps for the DDPM [24] denoising process following [53]. During training, the minimum and maximum length of motion sequences are set to 40 and 300 for both datasets. We insert an additional reference end latent at the end of each motion sequence to indicate the stop of generation. For Two-Forward strategy, a cosine scheduler is employed to control the ratio of replaced motion tokens, which can be formulated as: $\gamma_{t}=\frac{1}{2}(1-\cos(\frac{\pi t}{T}))$, where $t$ is current iteration step and $T$ is the total number of iterations. When $t=0$, $\gamma_{t}=0$, indicating that no generated motion tokens in the first forward pass are replaced, thus relying on the ground-truth motion tokens only. When $t=T$, $\gamma_{t}=1$, indicating that all generated motion tokens in the first forward pass are replaced, thus relying on the generated motion tokens only. We use the same optimizer as the Causal TAE and a batch size of 256. The initial learning rate is 1e-4 after 10K warmup iterations and decay to 0 for another 90K iterations using cosine learning rate scheduler. Our experiments are conducted on NVIDIA A800 GPUs.
@@
FID $\downarrow$ MPJPE $\downarrow$ FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
Real motion - - 0.002 0.711 0.851 0.903 15.805 27.670
(12,512) 8.862 38.5 21.078 0.600 0.759 0.827 17.143 27.755
@@
(20,1024) 0.379 19.9 11.010 0.630 0.765 0.847 16.802 27.485
(20,1280) 0.429 20.11 16.465 0.557 0.705 0.774 17.680 27.490
(20,1792) 0.548 20.1 11.145 0.616 0.776 0.842 16.919 27.597
(20,2048) 0.690 20.7 11.910 0.625 0.782 0.844 16.785 27.542
Table 6: Comparison with baseline motion tokenizers on HumanML3D [17] test set. MPJPE is measured in millimeters. (16, 1024) indicates the latent dimension and hidden size of the Causal TAE.
@@ D Classifier-free guidance
AR. layers AR. heads AR. dim Diff. layers FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
8 8 512 2 14.336 0.598 0.747 0.802 16.983 27.787
8 8 512 3 13.764 0.602 0.758 0.819 16.972 27.742
@@
E Failure of Inverse Kinematics
Post-processing for 263-dimentional motion representation.  Most previous works [53, 19, 60] uses 263-dimentional motion representation [17]. The representation can be written as follows:
$x=\{\dot{r}^{x},\dot{r}^{z},r^{y},\dot{r}^{a},\,j^{p},\,j^{v},\,j^{r},c\},$ (9)
where the root is projected on the XZ-plane (ground plane), $(\dot{r}^{x},\dot{r}^{z}\in\mathbb{R})$ are root linear velocities on the XZ-plane, $r^{y}\in\mathbb{R}$ is the root height, $\dot{r}^{a}\in\mathbb{R}^{1}$ denotes root angular velocity along the Y-axis, $j^{p}\in\mathbb{R}^{3(K-1)}$, $j^{v}\in\mathbb{R}^{3K}$, and $j^{r}\in\mathbb{R}^{6(K-1)}$ are local joint positions, local velocities, and local rotations relative to the root, $K$ is the number of joints (including the root), and $c\in\mathbb{R}^{4}$ is the contact label. For SMPL characters, we have $K=22$ and we get $2+1+1+3\times 21+3\times 22+6\times 21+4=263$ dimensions. In the original implementation [17], the joint rotation is directly solved using Inverse Kinematics (IK) with relative joint positions. In such way, the joint loses twist rotation and directly applying the joint rotation to the character faces a lot of rotation error [14], as shown in Fig. 8. To overcome this issue, previous works [53, 19, 60] only uses the positions and employs SMPLify [5] to solve the real SMPL joint rotation. This process is time-consuming (around $60$ seconds for a $10$ seconds motion clip) and also introduces unnatural results like jittering head [53].
Most data in the HumanML3D [17] dataset comes from the AMASS [35] dataset. As the AMASS dataset provides the SMPL joint rotation, we slightly modify the motion representation by directly using the SMPL joint rotation and make it a 6D rotation for better learning. Consequently, we remove the slow post-processing step and easily drive the SMPL character with the generated rotations. The processing scripts to obtain our 272-dim motion representation are available at https://github.com/Li-xingXiao/272-dim-Motion-Representation.

### v3

@@ A Implementation Details
For the Causal TAE, both the encoder and decoder are based on the 1D causal ResNet blocks [23]. The temporal downsampling rate $l$ is set to 4 and all motion sequences are cropped to $N=64$ frames during training. We train the first 1900K iterations with a learning rate of 5e-5 and the remaining 100K iterations with a learning rate of 2.5e-6. We use the AdamW optimizer [38] with $\beta_{1},\beta_{2}]=[0.9,0.99]$ and a batch size of 128. We provide an ablation study on the hyperparameter $\lambda$ of root loss $L_{root}$ in Tab. 5. The latent dimension $d_{c}$ and hidden size are set to 16 and 1024, respectively. The latent dimension significantly impacts the compression rate, while the hidden size affects the model’s capacity. Both factors influence reconstruction and subsequent generation quality, requiring a careful trade-off between compression efficiency and generative performance. Ablation studies on the latent dimension and hidden size are provided in Tab. 6. To further improve the quality of the reconstructed motion, we add a linear layer after the embedded Gaussian distribution parameters as a latent adapter to get a lower-dimensional and more compact latent space for subsequent sampling, as proposed in [13].
For the Transformer inside the AR model, we use the architecture akin to LLaMA [66] with 12 layers, 12 attention heads and 768 hidden dimension. The ablation for different scales of the Transformer is provided in Tab. 7. Block size is set to 78 and we choose RoPE [63] as the positional encoding. For the diffusion head after Transformer, we use MLPs with 1792 hidden dimension and 9 layers. The output vectors of the Transformer serve as the condition of denoising via AdaLN [46]. We adopt a cosine noise schedule with 50 steps for the DDPM [27] denoising process following [64]. During training, the minimum and maximum length of motion sequences are set to 40 and 300 for both datasets. We insert an additional reference end latent at the end of each motion sequence to indicate the stop of generation. For Two-Forward strategy, a cosine scheduler is employed to control the ratio of replaced motion tokens, which can be formulated as: $\gamma_{t}=\frac{1}{2}(1-\cos(\frac{\pi t}{T}))$, where $t$ is current iteration step and $T$ is the total number of iterations. When $t=0$, $\gamma_{t}=0$, indicating that no generated motion tokens in the first forward pass are replaced, thus relying on the ground-truth motion tokens only. When $t=T$, $\gamma_{t}=1$, indicating that all generated motion tokens in the first forward pass are replaced, thus relying on the generated motion tokens only. We use the same optimizer as the Causal TAE and a batch size of 256. The initial learning rate is 1e-4 after 10K warmup iterations and decay to 0 for another 90K iterations using cosine learning rate scheduler. Our experiments are conducted on A800 GPUs.
@@
FID $\downarrow$ MPJPE $\downarrow$ FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ $M$-D $\downarrow$ Div $\rightarrow$
Real motion - - 0.002 0.702 0.864 0.914 15.151 27.492
(12,512) 8.862 38.5 21.078 0.600 0.759 0.827 17.143 27.456
@@
(20,1024) 0.379 19.9 11.814 0.630 0.765 0.847 16.802 27.485
(20,1280) 0.429 20.1 16.465 0.557 0.705 0.774 17.680 27.490
(20,1792) 0.548 20.1 11.845 0.616 0.776 0.842 16.919 27.392
(20,2048) 0.690 20.7 11.910 0.625 0.782 0.844 16.785 27.346
Table 6: Ablation Study of different Causal TAE architecture designs on HumanML3D [20] test set. Each generation model remains the same. MPJPE is measured in millimeters. (16, 1024) indicates the latent dimension and hidden size of the Causal TAE.
@@ D Classifier-free guidance
AR. layers AR. heads AR. dim Diff. layers FID $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ $M$-D $\downarrow$ Div $\rightarrow$
8 8 512 2 14.336 0.598 0.747 0.802 16.983 27.287
8 8 512 3 13.764 0.602 0.758 0.819 16.972 27.242
@@
E Failure of Inverse Kinematics
Post-processing for 263-dimensional motion representation.  Most previous works [64, 22, 76] uses 263-dimensional motion representation [20].
The representation can be written as follows:
$x=\{\dot{r}^{a},\dot{r}^{x},\dot{r}^{z},r^{y},\,j^{p},\,j^{v},\,j^{r},c\},$ (9)
where the root is projected on the XZ-plane (ground plane), $\dot{r}^{a}\in\mathbb{R}^{1}$ denotes root angular velocity along the Y-axis, $(\dot{r}^{x},\dot{r}^{z}\in\mathbb{R})$ are root linear velocities on the XZ-plane, $r^{y}\in\mathbb{R}$ is the root height, $j^{p}\in\mathbb{R}^{3(K-1)}$, $j^{v}\in\mathbb{R}^{3K}$, and $j^{r}\in\mathbb{R}^{6(K-1)}$ are local joint positions, local velocities, and local rotations relative to the root, $K$ is the number of joints (including the root), and $c\in\mathbb{R}^{4}$ is the contact label. For SMPL characters, we have $K=22$ and we get $2+1+1+3\times 21+3\times 22+6\times 21+4=263$ dimensions. In the original implementation [20], the joint rotation is directly solved using Inverse Kinematics (IK) with relative joint positions. In such way, the joint loses twist rotation and directly applying the joint rotation to the character faces a lot of rotation error [16], as shown in Fig. 8. To overcome this issue, previous works [64, 22, 76] only uses the positions and employs SMPLify [5] to solve the real SMPL joint rotation. This process is time-consuming (around $60$ seconds for a $10$ seconds motion clip) and also introduces unnatural results like jittering head [64]. Most data in the HumanML3D [20] dataset comes from the AMASS [41] dataset. As the AMASS dataset provides the SMPL joint rotation, we slightly modify the motion representation by directly using the SMPL joint rotation and make it a 6D rotation for better learning. Consequently, we remove the slow post-processing step and easily drive the SMPL character with the generated rotations. The processing scripts to obtain our 272-dim motion representation are available at https://github.com/Li-xingXiao/272-dim-Motion-Representation.
F Limitations and Future work
Limitations.  Despite its effectiveness, the streaming generation paradigm limits the applications of motion in-betweening and localized editing of intermediate tokens, as it inherently relies on unidirectional modeling. This limitation restricts flexibility in scenarios requiring fine-grained adjustments, such as seamlessly inserting new motions between existing frames or interactively refining motion details while preserving global coherence. Future work.  Future work could explore hybrid strategies that allow bidirectional refinement without compromising streaming generation. One potential way is to predict a set of future latents at each step, which could enable motion in-between and localized editing while preserving streaming manner.